OpenStudy (anonymous):

What values of x and y satisfy x + 2y = 6 and 2x + 3y = 4 simultaneously? A) x = 0 and y = 3 B) x = 2 and y = 0 C) x = -10 and y = 8 D) x = 8 and y = -1 @Salmon @kelliegirl33

OpenStudy (anonymous):

should be option 'c'

OpenStudy (anonymous):

Right! :) Thanks.

OpenStudy (anonymous):

x + 2y = 6 -->(-2) 2x + 3y = 4 -------------- -2x - 4y = -12 2x + 3y = 4 -------------- -y = - 8 y = 8 x + 2y = 6 x + 2(8) = 6 x + 16 = 6 x = 6 - 16 x = -10 x = -10 and y = 8............I agree..answer is C